What is the wholesale price of a CD player if the percent increase is 75% and the amount of increase is $30?
Answer by ikdeep(226)

% Increase in price = 75%
Amount increase in price = $30
Let the total amount after increase in pirce = $x
SO we get the equation as ....
75% of x = $ 30
75 /100 * x = 30
0.75 * x = 30

x = 30 / 0.75

x = $40
Therfore total amount after increase in pirce = $40
Now to calculate original or wholesale price .....
Total amount after increase in price = original price + increase in price

$40 = oirginal price + $30
40 - 30 = original price
or original or whole price = $10

Please feel free to revert back for any further queries.